I don't understand how to find the area of a segment of a circle.
For example the question says, Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius.

Answer by Alan3354(21581)

Find the area of the smaller segment whose chord is 8" long in a circle with an 8" radius.
--------------
Since all sides are 8", they form an equilateral triangle.
The area of the sector is pi*r^2*60/360
Subtract the area of the triangle, that leaves the sector.
---------------
Sector = pi*64*/6 = 32pi/3 sq inches
Area of triangle = 4*8*cos(30) = 16sqrt(3)
Area of segment = (32pi/3) - 16sqrt(3)